OpenStudy (anonymous):

A number is divided by 2. If 16 is added to the quotient you get 20. What is the number? 2 8 18 72

OpenStudy (math&ing001):

20-16 is the quotient, meaning 4*2 is the number.

OpenStudy (anonymous):

i dont get it

OpenStudy (math&ing001):

We got \(\frac{number}{2}+16 = 20 \) so \(\frac{number}{2}=20-16=4\) meaning \(number=4*2=8\) . I hope this is more clear.